At the park, I saw boys and dogs. I saw thirty two heads and 104 legs. How many were dogs and how many boys? Which equations did you use? I got 12 boys and 20 dogs by trial and error but I know that there must be some equations to solve. Please help. Thank you.
Answer by mananth(16949)

total 104 legs
32 heads
..
let there be x boys & y dogs
..
2x+4y=104..................1
x+y=32......................2
multiply equation 2 by -2 and add it to equation 1
-2x-2y=-64
add
4y-2y=104-64
2y=40
y=20 dogs
32-20= 12 boys
...
12*2+20*4=104